Circular No. 6756

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

NOVA IN THE LARGE MAGELLANIC CLOUD 1997 The MACHO collaboration (cf. IAUC 6312) reports their discovery

of an apparent nova in the Large Magellanic Cloud at α = 5^{h}04^{m}26^{s}.7, δ = −67^{o}38′38′′ (equinox 2000.0). The nova peaked in brightness sometime between consecutive observations on June 3.398 UT (prior to eruption) and June 16.395 (when the star was at least 8.5 mag brighter than on June 3, at V ∼ 13.5 or brighter). Currently, the nova is at V ∼ 17.9.

GRB 970828 S. Klose, J. Eisloeffel, and B. Stecklum, Thüringer Landessternwarte

Tautenburg, report: “The GRB error box was surveyed down to K′ = 20 with the Calar Alto 3.5-m telescope on Sept. 9.8 UT using the MAGIC nearinfrared camera (0′′.81 pixels). A comparison with R- and I-band images obtained with the prime-focus CCD camera of the Tautenburg Schmidt telescope on Aug. 30.9, 31.9, Sept. 1.9, and 6.9, and with R-band images taken by Stanek et al. (IAUC 6735) and by Howk et al. (3.5-m WIYN

telescope; cf. http://www.noao.edu/noao/grb/970828.html), shows that

all but one object visible in K′ within the revised ASCA error circle (IAUC 6732) have a counterpart in R and/or I, down to R = 22–23 and I = 20. This one object at α = 18^{h}08^{m}30^{s}.52, δ = +59^{o}18′49′′.4 (± 1′′, equinox 2000.0) is notably fainter. Its magnitudes are R = 23–24 (?) and K′ = 19.0 ± 0.4. Since no variable was found in the error box down to R = 24.5 (Odewahn et al., IAUC 6735), we suppose, however, that we have not detected the GRB afterglow and constrain its K′ magnitude 12 days after the burst to be > 20. Images are available via http://www.tls-

tautenburg.de/research/research.html.”

COMET C/1995 O1 (HALE-BOPP) Total visual magnitude estimates: Sept. 16.76 UT, 4.6 (D. A. J. Sear-

gent, Cowra, N.S.W., naked eye; during total lunar eclipse); 24.32, 4.8 (J. G. de S. Aguiar, Campinas, Brazil, 11×80 binoculars); 30.44, 5.1 (G. W. Kronk, Troy, IL, 20×80 binoculars); Oct. 4.32, 5.3 (R. Lourençon, Jundiai, SP, Brazil, 20×80 binoculars); 10.84, 5.4 (A. Pearce, Subiaco, W. Australia, 20×80 binoculars); 12.30, 5.2 (Aguiar); 13.84, 5.6 (Pearce).

1997 October 14  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Daniel W. E. Green